# Data Caricatures: On the Representation of African American Language in Pretraining Corpora

### v1

@@ Abstract
With a combination of quantitative experiments, human judgments, and qualitative analyses, we evaluate the quantity and quality of African American Language (AAL) representation in 12 predominantly English, open-source pretraining corpora. We specifically focus on the sources, variation, and naturalness of included AAL texts representing the AAL-speaking community. We find that AAL is underrepresented in all evaluated pretraining corpora compared to US demographics, constituting as little as 0.007% of documents. We also find that more than 25% of AAL texts in C4 may be inappropriate for LLMs to generate and reinforce harmful stereotypes. Finally, we find that most automated language, toxicity, and quality filters are more likely to conserve White Mainstream English (WME) texts over AAL in pretraining corpora.
@@
[…]
@@ 1 Introduction
Figure 1: We evaluate the quantity and quality of automatically identified AAL documents in open source pretraining corpora. We find AAL is underrepresented across corpora and many documents contain texts that are misrepresentative of naturalistic AAL or reinforce stereotypes.
Given the close relationship between language and culture Adilazuarda et al. (2024), the distributions of sources that texts are drawn from may also misrepresent the broader culture of communities when used to train models. For example, many websites publish hip hop lyrics which may be collected and incorporated into pretraining data from online sources. While hip hop is an important component of Black culture and identity Clay (2003) as well as that of other communities, other registers of AAL (e.g., natural dialogue) are needed to more comprehensively capture the rich and diverse facets of Black culture. Furthermore, hip hop lyrics are not independently representative of AAL as a whole; lyrics do not necessarily reflect natural speech and no explicit indication of genre or register is provided during pretraining. Prior studies validate this gap, finding that Black users perceive a lack of cultural knowledge in current language technologies; they often suspect that AAL speakers are overlooked in model development processes Brewer et al. (2023); Cunningham et al. (2024).
Therefore, in addition to the quantity of AAL that is represented, we argue that the quality of the representation of AAL is also important. [^2] Note, biases may also come from the technical representation of text in the form of text embeddings. This is distinct, however, from the representation that we explore in this work, which refers to how the language of AAL speakers is depicted in pretraining corpora. We define quality as the extent to which texts faithfully and authentically portray the linguistic patterns–and consequently culture–of native AAL speakers. Accordingly, low quality representations of AAL (e.g., texts written by non-native speakers) pose representational harms to AAL speakers Blodgett et al. (2020).
We focus on three guiding research questions:
@@ 3. RQ3)
We use mixed-methods studies to address these questions, incorporating quantitative experiments, human judgments, and qualitative examination of the documents in pretraining corpora. To investigate RQ1 (Section 4), we measure the prevalence of AAL in 12 open-source pretraining corpora and find that as few as 0.007% of pretraining corpora documents contain AAL. To investigate RQ2 (Section 5), we conduct automated and human evaluations of AAL texts in pretraining corpora and find that a significant number of documents are unrepresentative of naturalistic AAL [^3] By naturalistic AAL, we refer to the linguistic patterns in text that most closely resemble everyday, spontaneous speech. (e.g., perceived to be written by non-native speakers). Surprisingly, some texts resembling AAL are actually posted by corporate social media accounts. Finally, to investigate RQ3 (Section 6), we evaluate 16 quality filtering approaches and find that most filters are more likely to conserve WME texts compared to AAL, particularly for texts from social media or song lyrics.
@@ 2 Background and Related Work
AAL, Performativity, and Misrepresentation. Research involving AAL in the NLP community has relied on texts drawn from different sources, such as speech transcripts Farrington and Kendall (2021) and social media Blodgett et al. (2016). Across such sources, AAL use can reflect varying degrees of what has been termed performative speech, language practices which use the rhetorical style of AAL such as signifyin’ Mitchell-Kernan and Thomas (1972) on a speaker’s cultural background to construct or communicate Black identity. Because they primarily rely on language to communicate Black culture, performative register is especially prevalent in hip hop Alim (2006) or social media language Eisenstein (2013); Ilbury (2020). Recent work has also argued that the linguistic patterns of these sources may misrepresent the patterns of AAL-speakers’ when interacting with LLMs Kleiner et al. (2024). Alternatively, AAL can be misrepresented through the derogatory use of features Ronkin and Karn (1999), diffusion of linguistic markers Corradini (2024), and use of AAL-like language in marketing Roth-Gordon et al. (2020). In this work, we examine the quality of AAL representation by considering performativity and these potential misrepresentations.
@@
AAL Biases in Language Technologies. The most prominent documentation and mitigation of AAL biases in text-based language models has focused on toxicity detection (e.g., Sap et al. 2019a; Harris et al. 2022; Davidson et al. 2019; Cheng et al. 2022; Halevy et al. 2021). More recent work evaluates AAL biases in other contexts, including classification tasks through synthetic data Ziems et al. (2022); Dacon et al. (2022), extractive summarization Keswani and Celis (2021); Olabisi et al. (2022), LLMs’ stereotyping behaviors Fleisig et al. (2024); Hofmann et al. (2024), and generative language models Groenwold et al. (2020); Deas et al. (2023, 2024). Rather than evaluate downstream biases, we specifically examine AAL in pretraining data as a potential source of bias.
@@
Pretraining Data Quality. Hovy and Prabhumoye (2021) identify data as a fundamental source of bias, and prior work has attempted to trace various biases to training data (e.g., Ladhak et al. (2023); Feng et al. (2023); Dodge et al. (2021)). Recent work has also increasingly been concerned with measuring and selecting high quality data for pretraining LLMs (e.g., Wettig et al. (2024)). Such studies, however, use a variety of different measures of data quality. A majority of prior work (e.g., Li et al. 2024) primarily operationalizes quality through downstream model performance on broad benchmarks (e.g., MMLU Hendrycks et al. (2021) or HellaSwag Zellers et al. (2019)). Other work measures quality using sources that are typically considered "high-quality" (e.g., Wikipedia) in developing automated data filters (e.g., RedPajama, Computer (2023); The Pile, Gao et al. (2020)). Finally, relatively fewer studies have examined intrinsic notions of quality, such as educational value or lack of harmful content Wettig et al. (2024); Sachdeva et al. (2024); Gunasekar et al. (2023). With the aim of identifying potential causes of AAL biases, we measure quality through a variety of means, focusing on texts’ sources, representativeness of naturalistic AAL, and lack of harmful reflections on AAL-speakers.
@@ 3 Data Extraction
Dataset Example Models # Docs (sample) % CC Docs # AAL Docs (%)
@@
OpenWebText Gokaslan et al. (2019) SSDLM, RoBERTa 8M 0% 2,684 (.03%)
@@
Table 1: The 12 open-source pretraining datasets evaluated, including example models pretrained on each dataset, size in raw unique documents, and proportion composed of Common Crawl. indicates only non-copyrighted portions are available and included in analyses. Bolded corpora indicate that C4 is explicitly included in the corpus before further filtering. Given their size, the bottom 4 corpora are analyzed using a random sample of dataset shards; 99% confidence intervals on each sampled corpora’s estimate is included.
We examine the quantity and quality of AAL texts present in 12 predominantly English, open-source pretraining corpora. All corpora are listed in Table 1. In particular, we consider FineWeb-Edu (italicized in Table 1) as a baseline, given that it strictly prioritizes highly educational content (e.g., Wikipedia) as opposed to diverse everyday language use. Following prior work’s documentation of C4 Dodge et al. (2021) and because many corpora predominantly rely on Common Crawl texts, we focus our human judgments and AAL feature analyses on variants of the C4 corpus Raffel et al. (2020).
@@ 3.1 Extracting AAL Subsets
To identify AAL texts, we follow Dodge et al.’s (2021) analysis of C4 as well as other related work (e.g., Xia et al. (2020); Sap et al. (2019b); Davidson et al. (2019)) and use a mixed-membership demographic-alignment classifier validated with common linguistic features of AAL Blodgett et al. (2016). From eight full corpora, we extract all texts where AAL is the most likely classification (additional corpora details are included in Subsection A.1). For four exceedingly large corpora with more than 3 billion documents–Dolmino/Olmo-mix OLMo et al. (2024), DCLM-baseline Li et al. (2024), FineWeb Penedo et al. (2024), and RedPajama-v2 Weber et al. (2024)–we analyze a 250 GB sample of each corpus. Additional extraction methodology and corpus details are included in Subsection A.2 and Appendix B.  [^4] While corpora like RedPajama-2 are intended to enable experimentation with automatic filtering rather than to be used as is for pretraining, we refer to all corpora as ”pretraining corpora” for simplicity. To account for AAL features that may be present within documents that largely reflect WME, we additionally extract a more conservative subset of AAL texts using a threshold of 0.3 for these analyses.  [^5] While prior work largely considers a threshold of 0.8 (e.g., Xia et al. (2020)), we choose a threshold of 0.3 to calculate a more conservative estimate of feature prevalence while maintaining a manageable corpus size for analysis. Additional discussion is included in Appendix C.
We use the same block list of terms used in the original C4 corpus Raffel et al. (2020) to identify texts that would be filtered from the corpus before pretraining. This filtering yields two distinct variants: C4.en.noBlocklist which does not employ the block list, and C4.en which lacks block list-identified texts. With these two variants, we conduct a fine-grained evaluation of the block list’s impacts on AAL representation, prevalence of features, and human judgments of AAL texts.
@@ 4.1 Methods
We recruit three annotators to conduct the human judgments of the sampled texts; all annotators are self-reported native AAL-speakers. Annotators also currently study or have previously studied linguistics or computational linguistics. Annotators are provided a text from either the AAL subset of C4.en.noBlocklist or C4.en and asked to label each text first on two dimensions drawn from prior work Deas et al. (2023): Human-Likeness asks whether the text appears authored by a human and Linguistic Match asks whether there are identifiable features of AAL in the given text. Both dimensions are judged on 4-point Likert scales (higher representing more human like and more identifiable features). Among all three annotators, we collect judgments for 1,054 texts. Annotators exhibit moderate to substantial agreement for binarized Human-Likeness and Linguistic Match dimensions ($\kappa=0.581$ and $\kappa=0.747$ respectively). Additional sampling, annotation, and interface details are included in Appendix D.
Automated Feature Extraction. To conduct a more fine-grained analysis, we estimate the distribution of morphosyntactic AAL features. Features are identified using the CGEdit model Masis et al. (2022), a classifier developed with a human-in-the-loop framework. The classifier considers 17 AAL morphosyntactic constructions, including features like habitual be (e.g., "He be driving") and copula deletion (e.g., "He $\emptyset$ at home"). See Masis et al. 2022 and Appendix E for a detailed list of features. As the classifier was shown to be reliable for examining distributions in large collections of text, we restrict analysis of features to trends in large data subsets rather than individual texts.
@@ 4.2 Results
AAL Frequency. Table 1 includes the proportion of documents labeled as AAL in each corpus evaluated. Language understanding benchmarks drive pretraining data curation choices and are drawn from sources such as descriptive video captions (e.g., HellaSwag Zellers et al. (2019), Swag Zellers et al. (2018)) or academic exams (e.g., MMLU Hendrycks et al. (2021)). Therefore, we expect that the analyzed corpora contain few documents with AAL. In line with this intuition, we find that AAL is extremely underrepresented across corpora used and intended for LLM pretraining, included in as few .007% of documents (i.e. RedPajama). As expected, the FineWeb-Edu corpus, which employs among the strictest filtering approaches, exhibits the lowest percentage of documents containing AAL (.0009%), while RedPajama-v2, which employs relatively limited filtering, exhibits the highest percentage (.18%). Due to this trend and those found in prior work, we investigate filtering approaches further in Section 6.
AAL Prob. C4 # Docs % AAL Judgments
@@
To verify the automated analyses, Table 2 presents the proportion of texts labeled by annotators as containing identifiable AAL features within different ranges of AAL posterior probability as output by the demographic alignment classifier. Annotators indicate that a significant proportion of the texts do not contain features of AAL (55.3% in the $0.5\leq p\leq 0.6$ range). Many of these documents instead contain lists of search terms and artist names that resemble lexical markers of AAL (e.g., Lil Wayne) rather than naturalistic AAL.
AAL Feature Frequency. Figure 2 presents the frequency of each automatically detected AAL feature in C4.en.noBlocklist and C4.en. Supporting Dodge et al.’s (2021) finding that the block list disproportionately filters AAL texts, we see that filtering lowers the frequency of all AAL features. Some of the most frequent features in C4.en.noBlocklist–zero copula (ZC) and multiple negatives (MN)–are among the most frequent documented features of naturalistic AAL. Other features such as negative auxiliary inversion (NAI) and zero plurals (ZP), however, have extremely low representation despite being relatively common features among both urban and rural AAL-speakers Kortmann et al. (2020).
While some features are represented in frequencies that generally reflect naturalistic AAL, Figure 2 also illustrates how these frequencies are impacted by filtering. While the most frequent features are generally shared among both subsets, the order of most frequent features diverges. While Zero Copula (ZC; e.g., he $\emptyset$ running) is detected most often in C4.en.noBlocklist, it is far less represented in C4.en. This alteration suggests that filtering may impact AAL-speakers differently depending on region, class, and other speaker characteristics (e.g., Southern or working class; Wolfram and Kohn 2015). Not only does C4 filtering disproportionately remove AAL texts, but filtering also systematically alters the distribution of AAL features.

### v2

@@ Abstract
With a combination of quantitative experiments, human judgments, and qualitative analyses, we evaluate the quantity and quality of African American Language (AAL) representation in 12 predominantly English, open-source pretraining corpora. We specifically focus on the sources, variation, and naturalness of included AAL texts representing the AAL-speaking community. We find that AAL is underrepresented in all evaluated pretraining corpora compared to US demographics, constituting as few as 0.007% and at most 0.18% of documents. We also find that more than 25% of AAL texts in C4 may be perceived as inappropriate for LLMs to generate and to reinforce harmful stereotypes. Finally, we find that most automated filters are more likely to conserve White Mainstream English (WME) texts over AAL in pretraining corpora.  [^1] We make our code available at https://github.com/NickDeas/DataCaricatures
@@
Nicholas Deas*1, Blake Vente*1, Amith Ananthram1, Jessica A. Grieser2, Desmond Patton3, Shana Kleiner3, James Shepard4, Kathleen McKeown1 1 Columbia University, Department of Computer Science 2 University of Michigan, Department of Linguistics 3 University of Pennsylvania, School of Social Policy and Practice, Annenberg School for Communications 4 University of Tennessee, Knoxville, Department of English {ndeas, kathy}@cs.columbia.edu  rv2459@columbia.edu
@@ 1 Introduction
Figure 1: We evaluate the quantity and quality of automatically identified AAL documents in open source pretraining corpora. We find AAL is underrepresented across corpora and many documents contain texts that are misrepresentative of naturalistic AAL.
Given the close relationship between language and culture Adilazuarda et al. (2024), the distributions of sources that texts are drawn from may also misrepresent the broader culture of communities when used to train models. For example, many websites publish hip hop lyrics which may be collected and incorporated into pretraining data. While hip hop is an important component of Black culture and identity Clay (2003) as well as that of other communities, other registers of AAL (e.g., natural dialogue) are needed to more comprehensively capture the rich and diverse facets of Black culture. Furthermore, hip hop lyrics are not independently representative of AAL as a whole; lyrics do not necessarily reflect natural speech and no explicit indication of genre or register is provided during pretraining. Prior studies validate this gap, finding that Black users perceive a lack of cultural knowledge in current language technologies; they often suspect that AAL speakers are overlooked in model development processes Brewer et al. (2023); Cunningham et al. (2024).
Therefore, in addition to the quantity of AAL that is represented, we argue that the quality of its representation is also important. [^3] Note, biases may also come from the technical representation of text in the form of text embeddings. This is distinct, however, from the representation that we explore in this work, which refers to how the language of AAL speakers is depicted in pretraining corpora. We define quality as the extent to which texts authentically portray the linguistic patterns–and consequently culture–of native AAL speakers. Accordingly, low quality representations of AAL (e.g., texts written by non-native speakers) pose representational harms to AAL speakers Blodgett et al. (2020). We focus on three guiding research questions:
@@ 3. RQ3)
We use mixed-methods studies to address these questions, incorporating quantitative experiments, human judgments, and qualitative examination of the documents in pretraining corpora. To investigate RQ1 (Section 4), we measure the prevalence of AAL in 12 open-source pretraining corpora and find that 0.007% to 0.18% of documents among pretraining corpora we evaluate contain AAL. In contrast, 80% of African Americans ($\sim$10of Americans) are AAL speakers Green (2002). To investigate RQ2 (Section 5), we conduct automated and human evaluations of AAL texts in pretraining corpora and find that a substantial number of documents are unrepresentative of naturalistic AAL [^4] By naturalistic AAL, we refer to the linguistic patterns in text that most closely resemble everyday, spontaneous speech. (e.g., perceived to be written by non-native speakers). Surprisingly, some texts resembling AAL are actually posted by corporate social media accounts. Finally, to investigate RQ3 (Section 6), we evaluate 16 automated filtering approaches and find that most filters are more likely to conserve WME texts compared to AAL, particularly for texts from social media or song lyrics.
@@ 2 Background and Related Work
AAL, Performativity, and Misrepresentation. Research involving AAL in the NLP community has relied on texts drawn from different sources, such as speech transcripts Farrington and Kendall (2021) and social media Blodgett et al. (2016). Across such sources, AAL use can reflect varying degrees of what has been termed performative speech, language practices which use the rhetorical style of AAL such as signifyin’ Mitchell-Kernan and Thomas (1972) on a speaker’s cultural background to construct or communicate Black identity. Because they primarily rely on language to communicate Black culture, performative register is especially prevalent in hip hop Alim (2006) or social media language Eisenstein (2013); Ilbury (2020). Recent work has also argued that the linguistic patterns of these sources may misrepresent the patterns of AAL speakers’ when interacting with LLMs Kleiner et al. (2024). Alternatively, AAL can be misrepresented through the derogatory use of features Ronkin and Karn (1999), diffusion of linguistic markers Corradini (2024), and use of AAL-like language in marketing Roth-Gordon et al. (2020). In this work, we examine the quality of AAL representation by considering performativity and these potential misrepresentations.
@@
AAL Biases in Language Technologies. The most prominent documentation and mitigation of AAL biases in text-based language models has focused on toxicity detection (e.g., Sap et al. 2019a; Harris et al. 2022; Davidson et al. 2019; Cheng et al. 2022; Halevy et al. 2021). More recent work evaluates AAL biases in other contexts, including classification tasks through synthetic data Ziems et al. (2022); Dacon et al. (2022), summarization Keswani and Celis (2021); Olabisi et al. (2022), stereotyping behaviors Fleisig et al. (2024); Hofmann et al. (2024), generative language models Groenwold et al. (2020); Deas et al. (2023, 2024), and reward models Mire et al. (2025). In contrast, we specifically study AAL in pretraining data as a potential source of bias.
@@
Pretraining Data Quality. Hovy and Prabhumoye (2021) identify data as a fundamental source of bias, and prior work has attempted to trace various biases to training data (e.g., Ladhak et al. (2023); Feng et al. (2023); Dodge et al. (2021)). Recent work has also increasingly been concerned with measuring and selecting high quality data for pretraining LLMs (e.g., Wettig et al. (2024)). Such studies, however, use a variety of different measures of data quality. A majority of prior work (e.g., Li et al. 2024) primarily operationalizes quality through downstream model performance on broad benchmarks (e.g., MMLU Hendrycks et al. (2021) or HellaSwag Zellers et al. (2019)). Other work measures quality using sources that are typically considered "high-quality" (e.g., Wikipedia) in developing automated data filters (e.g., RedPajama, Computer (2023); The Pile, Gao et al. (2020)). Finally, relatively fewer studies have examined intrinsic notions of quality, such as educational value or lack of harmful content Wettig et al. (2024); Sachdeva et al. (2024); Gunasekar et al. (2023). With the aim of identifying potential causes of AAL biases, we measure quality through a variety of means, focusing on texts’ sources, representativeness of naturalistic AAL, and lack of harmful reflections on AAL speakers.
@@ 3 Data Extraction
Dataset Example Models # Docs (sample) % CommonCrawl Docs # AAL Docs (%)
@@
OpenWebText Gokaslan et al. (2019) SSDLM, RoBERTa 8M 0% 858 (.01%)
@@
Table 1: The 12 open-source pretraining datasets evaluated, including example models pretrained on each dataset, size in raw unique documents, and proportion composed of Common Crawl. indicates only non-copyrighted portions are available and included in analyses. Bolded corpora indicate that C4 is explicitly included in the corpus before further filtering. Given their size, the bottom 4 corpora are analyzed using a random sample of dataset shards; 99% confidence intervals on each sampled corpus’ estimate is included.
We examine the quantity and quality of AAL texts present in 12 predominantly English, open-source pretraining corpora. All corpora are listed in Table 1. We consider FineWeb-Edu (italicized in Table 1) as a baseline, given that it strictly prioritizes highly educational content (e.g., Wikipedia) over diverse everyday language use. Following prior work’s documentation of C4 Dodge et al. (2021) and because many corpora predominantly rely on Common Crawl texts, we focus our human judgments and AAL feature analyses on variants of the C4 corpus Raffel et al. (2020).
@@ 3.1 Extracting AAL Subsets
To identify AAL texts, we follow Dodge et al.’s (2021) analysis of C4 as well as other related work (e.g., Xia et al. (2020); Sap et al. (2019b); Davidson et al. (2019)) and use a mixed-membership demographic-alignment classifier validated with common linguistic features of AAL Blodgett et al. (2016). From eight full corpora, we extract all texts where AAL is the most likely classification (additional corpora details are included in Subsection A.1). For four exceedingly large corpora (each with more than 3 billion documents)–Dolmino (Olmo-mix) OLMo et al. (2024), DCLM-baseline Li et al. (2024), FineWeb Penedo et al. (2024), and RedPajama-v2 Weber et al. (2024)–we analyze a 250 GB sample of each corpus. Additional extraction methodology and corpus details are included in Appendices A and B [^5] While corpora like RedPajama-2 are intended to enable experimentation with automatic filtering rather than to be used as is for pretraining, we refer to all corpora as ”pretraining corpora” for simplicity. To account for AAL features that may be present within documents that largely reflect WME, we additionally extract a more conservative subset of texts using a threshold of 0.3 for further analyses.  [^6] While prior work largely considers a threshold of 0.8 (e.g., Xia et al. (2020)), we choose a threshold of 0.3 to calculate a more conservative estimate of feature prevalence while maintaining a manageable corpus size for analysis. Additional discussion is included in Appendix C.
We use the same block list of terms used in the original C4 corpus Raffel et al. (2020) to identify texts that would be filtered from the corpus before pretraining. This filtering yields two distinct variants: C4.en.noBlocklist which does not employ the block list, and C4.en which lacks block list-identified texts Dodge et al. (2021). With these two variants, we conduct a fine-grained evaluation of the block list’s impacts on AAL representation, prevalence of features, and human judgments of AAL texts.
@@ 4 RQ1: How Much is AAL Represented?
We first aim to quantify the prevalence of AAL in pretraining corpora. We conduct our analysis through both human judgments of documents as well as automated approaches. In this section, we first measure the proportion of AAL documents in each corpus and the count of overlapping AAL documents between corpora. We then consider the representation of individual AAL features.
@@ 4.1 Methods
We recruit three annotators to conduct the human judgments of the sampled texts; all annotators are self-reported native AAL speakers. Annotators also currently study or have previously studied linguistics or computational linguistics. Annotators are provided a text from either the AAL subset of C4.en.noBlocklist or C4.en and asked to label each text first on two dimensions drawn from prior work Deas et al. (2023): Human-Likeness asks whether the text appears authored by a human and Linguistic Match asks whether there are identifiable features of AAL. Both dimensions are judged on 4-point Likert scales (higher representing more human-like and more reflective of AAL). We collect judgments for a total of 1,054 texts. Annotators exhibit moderate to substantial agreement for binarized Human-Likeness and Linguistic Match dimensions ($\kappa=0.581$ and $\kappa=0.747$ respectively). Additional sampling, annotation, and interface details are included in Appendix D.
Automated Feature Extraction. To conduct a more fine-grained analysis, we estimate the distribution of morphosyntactic AAL features with automated methods. Features are identified using the CGEdit model Masis et al. (2022), a classifier developed with a human-in-the-loop framework. The classifier considers 17 AAL morphosyntactic constructions, including features like habitual be (e.g., "He be driving") and copula deletion (e.g., "He $\emptyset$ at home"). See Masis et al. 2022 and Appendix E for a detailed list of features. As the classifier was shown to be reliable for examining distributions in large collections of text, we restrict analysis of features to trends in large data subsets rather than individual texts.
@@ 4.2 Results
AAL Frequency. Table 1 includes the proportion of documents labeled as AAL in each corpus evaluated. Language understanding benchmarks, in part, drive pretraining data curation choices and are drawn from sources such as descriptive video captions (e.g., HellaSwag Zellers et al. (2019), Swag Zellers et al. (2018)) or academic exams (e.g., MMLU Hendrycks et al. (2021)). Therefore, we expect that the analyzed corpora contain few documents with AAL. In line with this intuition, we find that AAL is extremely underrepresented across corpora used and intended for LLM pretraining, included in as few .007% of documents (i.e., RedPajama). As expected, the FineWeb-Edu corpus, which employs among the strictest filtering approaches, exhibits the lowest percentage of documents containing AAL (.0009%), while RedPajama-v2, which employs relatively limited filtering, exhibits the highest percentage (.18%). Due to this trend and those found in prior work, we investigate filtering approaches further in Section 6.
AAL Prob. # C4 Docs % AAL Judgments
@@
To verify the automated analyses, Table 2 presents the proportion of texts labeled by annotators as containing identifiable AAL features within different ranges of AAL posterior probability as output by the demographic alignment classifier. Annotators indicate that a significant proportion of the texts do not contain features of AAL (55.3% in the $0.5\leq p\leq 0.6$ range). We note that while the percentage of AAL judgments drops as AAL probability increases, this may be due to the classifier’s sensitivity to spurious token-level features, such as abbreviations in search terms and artist names (e.g., "Lil Wayne") that resemble lexical markers of AAL. To account for this, in our AAL feature and hip hop analyses, we consider all documents exceeding an AAL probability threshold of 0.3. See Appendix F for further discussion and additional examples of classifier predictions.
AAL Document Overlap. Figure 2 presents the count of overlapping AAL documents among the corpora analyzed. While some corpora explicitly include others (e.g., Dolma is a superset of C4), these results suggest that there is substantial overlap in the AAL documents included for each corpus. Among all AAL documents considered, 17% are duplicated in at least one other corpus. Therefore, not only is AAL underrepresented, but there is also a lack of diversity across corpora.
Figure 2: Counts of overlapping AAL documents among sets of corpora (indicated by checkmarks below the plot). Sizes of the AAL subsets for individual corpora are noted in parentheses. Overlapping sets smaller than 500 documents are not shown.
AAL Feature Frequency. Figure 3 presents the frequency of each automatically detected AAL feature in C4.en.noBlocklist and C4.en. Supporting Dodge et al.’s (2021) finding that the block list disproportionately filters AAL texts, we see that filtering lowers the frequency of all AAL features. Some of the most frequent features in C4.en.noBlocklist–zero copula (ZC) and multiple negatives (MN)–are among the most frequent documented features of naturalistic AAL. Other features such as negative auxiliary inversion (NAI) and zero plurals (ZP), however, have extremely low representation despite being relatively common features among both urban and rural AAL speakers Kortmann et al. (2020).
While some features are represented in frequencies that generally reflect naturalistic AAL, Figure 3 also illustrates how these frequencies are impacted by filtering. While the most frequent features are generally shared among both subsets, the order of most frequent features diverges. While Zero Copula (ZC; e.g., he $\emptyset$ running) is detected most often in C4.en.noBlocklist, it is far less represented in C4.en. This alteration suggests that filtering may behave differently on AAL texts depending on the speaker’s region, class, and other characteristics (e.g., Southern or working class; Wolfram and Kohn 2015). Not only does C4 filtering disproportionately remove AAL texts, but filtering also systematically alters the distribution of AAL features. Feature analyses of the remaining corpora are included in Appendix G, where we also observe substantial variation in feature distributions among different corpora.